OpenStudy (anonymous):

Use the formula for Interest: P = 500 r = 6.6% Y = 2 A = ?

OpenStudy (anonymous):

@CollegeCountryGuy we can;t give u exact answers

OpenStudy (anonymous):

DO u need explanations on the formula though

OpenStudy (anonymous):

Also u should tell us what y and x is standing for. Everyone has their own varuables for these things.

OpenStudy (bradely):

you need to compound or simple interest

OpenStudy (anonymous):

I understand just work through it with me @leejaeho2002

OpenStudy (anonymous):

college ok, but i need to know what y and a mean is it time, etc.

OpenStudy (bradely):

how many times of interest

OpenStudy (anonymous):

y is years and a is amount

OpenStudy (bradely):

A=P(1+(r/100))^n P= 500 r=6.6 n=2 A=500(1+(6.6/100))^2 =500(1.066)^2 = 568.178

OpenStudy (anonymous):

WEll the formula is p=a times y times r P = 500 r = 6.6% Y = 2 A = ?

OpenStudy (anonymous):

So we would have 500=a times 2 times 0.066

OpenStudy (bradely):

Further any assistance asked in www.mathskey.com

OpenStudy (anonymous):

0.132a=500